# Motivic interpretations for iterated integrals on some specific algebraic curves

### v2

@@ 3.2. Explicit formula
$\displaystyle\mathrm{I}_{h}(\omega_{4}^{2m}\omega_{0})$ $\displaystyle=\frac{1}{3^{2m}}\frac{1}{(2m)!}\left(\frac{\pi}{3}\right)^{2m}\frac{1}{2}\log 3+(-1)^{-m}\frac{1}{12^{m}}\zeta(2m+1)$ (92)
@@
$\displaystyle=\frac{1}{\sqrt{3}^{k-1}}\frac{1}{(k-1)!}\left\{\left(\frac{\pi}{3}\right)^{k-1}\log\frac{\sqrt{3}}{2}+(-1)^{k}(k-2)!(\mathrm{Li}_{k}(1)+\mathrm{Li}_{k}(-1))\right.$ (103)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\sum_{j=2}^{k}(-1)^{j-1}\frac{(k-1)!}{(k-j)!}\left(\frac{\pi\sqrt{-1}}{3}\right)^{k-j}\left(\mathrm{Li}_{j}(\xi_{6})-\frac{\pi\sqrt{-1}}{2}\frac{1}{(j-1)!}\left(\frac{\pi\sqrt{-1}}{3}\right)^{j-1}+\mathrm{Li}_{j}(-\xi_{6})\right)$ (104)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left.-\frac{\log 2}{k-1}\left(\frac{\pi\sqrt{-1}}{3}\right)^{k-1}+\frac{1}{k-1}\frac{\pi\sqrt{-1}}{2}\left(\frac{\pi\sqrt{-1}}{3}\right)^{k-1}-\frac{1}{k}\left(\frac{\pi\sqrt{-1}}{3}\right)^{k}\right\}.$ (105)
@@
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1}{\sqrt{3}^{k-2}}\sum_{j=2\text{: even}}^{k}\frac{(-1)^{\frac{2-j}{2}}}{(k-j)!}\left(\frac{\pi}{3}\right)^{k-j}2^{-j}L(j,\chi_{-3})-\frac{1}{\sqrt{3}^{k-1}}\sum_{j=3\text{: odd}}^{k}\frac{(-1)^{\frac{1-j}{2}}}{(k-j)!}\left(\frac{\pi}{3}\right)^{k-j}2^{-j}(3^{1-j}-1)\zeta(j)$ (110)
and the equations 91 and 94 are shown by setting $k=2m$ and $k=2m+1$. The latter claims are also shown in the same way as the proof of Proposition 3.5. ∎
@@ 4.1. Motivic iterated integrals
Let $\overline{\mathbb{Q}}$ be an algebraic closure of $\mathbb{Q}$ and $F(\subseteq\overline{\mathbb{Q}})$ be a number field. Then, then there are a graded commutative Hopf algebra $\mathcal{A}_{F}$ and a graded commutative $\mathcal{A}_{F}$-comodule $\mathcal{H}_{F}$. We investigate the motivic structure of periods in these algebras. Here, we review some properties of them and define motivic iterated integrals.
@@
Let $\mathcal{A}_{F}$ be the fundamental Hopf algebra of the abelian category $\mathrm{MTM}_{F}$ of mixed Tate motives over $F$. Put $\mathcal{H}_{F}:=\mathcal{A}_{F}\otimes_{\mathbb{Q}}\mathbb{Q}[\tau]$ where $\tau$ is a formal element. Then, $\mathcal{A}_{F}$ and $\mathcal{H}_{F}$ satisfy following conditions.
@@
The period map $\mathrm{per}:\mathcal{H}_{F}\rightarrow\mathbb{C}$ is conjecturally injective (called period conjecture, see [5, p.342, Conjecture 4.118]). This conjecture suggests that information about the structure of the periods is not lose by the motivic lifting.
@@
    $\mathrm{I}^{\mathfrak{a}}(p;a_{1},\dots,a_{k};q):=\rho(\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;a_{1},\dots,a_{k};q))\in\mathcal{A}_{N}^{(k)}.$
@@
    (path connection): For smooth paths $\gamma_{1},\gamma_{2}$ on $\mathbb{C}$,
@@
From here, we investigate the structures of $\mathcal{A}_{F}$ and $\mathcal{H}_{F}$ more precisely. For a $\mathbb{Q}$-subspace $R\subseteq F^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$, let $\mathcal{A}_{F,R}$ be a fundamental Hopf algebra of $R$. Then, it can be embedded into $\mathcal{A}_{F}$ (see [8, p.279, Appendix A.4]). We put $\mathcal{H}_{F,R}:=\mathcal{A}_{F,R}\otimes_{\mathbb{Q}}\mathbb{Q}[\tau]$.
@@
Here, $R_{N}:=\langle 1-\xi_{N}^{r}\leavevmode\nobreak\ |\leavevmode\nobreak\ r=1,\dots,N-1\rangle_{\mathbb{Q}}\subseteq\mathbb{Q}(\xi_{N})^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$.
@@ 4.2. Specific cases
In this subsection, we review some properties of the structure of $\mathcal{H}_{F,R}$ for some specific cases used in Section 5. We consider the two cases: (I) $(F,R)=(\mathbb{Q}(\xi_{4}),\langle 2\rangle_{\mathbb{Q}})$, which relates to multiple $L$-values of level 4; (II) $(F,R)=(\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{\mathbb{Q}})$, which relates to multiple $L$-values of level 6.
@@
    The $\mathbb{Q}$-linear space $\mathcal{H}_{\mathbb{Q}(\xi_{4}),\langle 2\rangle_{\mathbb{Q}}}^{(k)}$ is spanned by motivic iterated integrals formed
@@
    The $\mathbb{Q}$-linear space $\mathcal{H}_{\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{\mathbb{Q}}}^{(k)}$ is spanned by motivic iterated integrals formed
@@
    For $k\in\mathbb{Z}_{\geq{1}}$, the elements formed $\mathrm{I}^{\mathfrak{a}}(0;\delta,\{0\}^{k-1};1)$ $(\delta\in\mu_{4})$ are subject only to the following relations in $\mathcal{A}_{\mathbb{Q}(\xi_{4}),\langle 2\rangle_{\mathbb{Q}}}$:
@@
    For $k\in\mathbb{Z}_{\geq{1}}$, the elements formed $\mathrm{I}^{\mathfrak{a}}(0;\delta,\{0\}^{k-1};1)$ $(\delta\in\mu_{6})$ are subject only to the following relations in $\mathcal{A}_{\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{\mathbb{Q}}}$:
@@
$\displaystyle\mathrm{Ker}\Delta^{\prime}\cap\mathcal{H}_{\mathbb{Q}(\xi_{4}),\langle 2\rangle_{\mathbb{Q}}}^{(k)}=\mathbb{Q}\tau^{k}\oplus\mathbb{Q}\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{4},\{0\}^{k-1};1);$ (129)
$\displaystyle\mathrm{Ker}\Delta^{\prime}\cap\mathcal{H}_{\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{\mathbb{Q}}}^{(k)}=\mathbb{Q}\tau^{k}\oplus\mathbb{Q}\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{6},\{0\}^{k-1};1).$ (130)
@@ 5.1. Base extension
    The integral path is $\alpha:[0,1]\rightarrow\mathbb{P}^{1}(\mathbb{C})$, $\alpha(t)=1-\sqrt{\frac{1-t}{1+t}}$ and it satisfies $\alpha([0,1])=[0,1]$. Then, each $\mathrm{I}_{g}(\eta_{1}\cdots\eta_{k})$ can be written by a $\mathbb{Q}(\xi_{4})$-linear combination of iterated integrals on $\mathbb{P}^{1}(\mathbb{C})\setminus\{0,1,\pm\sqrt{1},\infty\}$ formed
@@
    with $a_{j}\in F_{g}:=\mathbb{Q}(\xi_{4})$ for $j=1,\dots,k$. Also, for each of the terms formed as the equation 131, it holds $\widetilde{\mathrm{I}}(a_{i_{1}},a_{i_{2}},a_{i_{3}})\in\langle 2\rangle_{\mathbb{Q}}=:R_{g}\subseteq F_{g}^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ if $1\displaystyle 0\leq i_{1}<i_{2}<i_{3}\leq k+$ 1.
@@
    $h(X,Y,Z)=X^{2}+XY+Y^{2}-Z^{2}$: for each element of $B_{h}=\left\{\omega_{0}:=\frac{\leavevmode\nobreak\ dx}{x},\omega_{1}:=\frac{\leavevmode\nobreak\ dx}{1-x},\omega_{4}:=\frac{\leavevmode\nobreak\ dx}{x+2y},\right.$
    $\left.\omega_{5}:=\frac{\leavevmode\nobreak\ dx}{x(x+2y)},\omega_{6}:=\frac{\leavevmode\nobreak\ dx}{(1-x)(x+2y)}\right\}$, the pullback along $\phi_{h}$ is
@@
    The integral path is $\alpha:[0,1]\rightarrow\mathbb{P}^{1}(\mathbb{C})$, $\alpha(t)=\frac{t-2+\sqrt{4-3t^{2}}}{2(1-t)}$ and it satisfies $[0,1]$. Then, each $\mathrm{I}_{h}(\eta_{1}\cdots\eta_{k})$ can be written by a $\mathbb{Q}(\xi_{6})$-linear combination of iterated integrals on $\mathbb{P}^{1}(\mathbb{C})\setminus\{0,1,-2,\xi_{3}^{\pm 1},\infty\}$ formed
@@
    with $a_{j}\in F_{h}:=\mathbb{Q}(\xi_{6})$ for $j=1,\dots,k$. Also, for each of the terms formed as the equation 137, it holds $\widetilde{\mathrm{I}}(a_{i_{1}},a_{i_{2}},a_{i_{3}})\in\langle 2,3\rangle_{\mathbb{Q}}=:R_{h}\subseteq F_{h}^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ if $1\displaystyle 0\leq i_{1}<i_{2}<i_{3}\leq k+$ 1.
By the above discussion, for each $f\in\{g,h\}$ and $(\eta_{1},\dots,\eta_{k})\in B_{f}^{(k)}$, $\mathrm{I}_{f}(\eta_{1}\cdots\eta_{k})$ can be written as an $F_{f}$-linear combination of iterated integrals formed
@@
with $a_{j}\in F_{f}$ for $j=1,\dots,k$. We review that the $\mathbb{Q}$-linear combination of the equation 138 has a motivic interpretation given by motivic iterated integrals in $\mathcal{H}_{F_{f},R_{f}}$. Then, we consider the base extension of $\mathcal{H}_{F_{f},R_{f}}$ to give motivic interpretations of iterated integrals on $X_{f}$.
@@
For a number field $F\supseteq\mathbb{Q}$ and $R\subseteq F^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$, we extend the following notations to $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$. For $u\otimes c,u_{1}\otimes c_{1},u_{2}\otimes c_{2}\in\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$,
@@
As same as [14, p.19, Proposition 4.4.2], if the period conjecture is true, (i.e. the period map $\mathrm{per}:\mathcal{H}_{F,R}\rightarrow\mathbb{C}$ is injective), then its extension $\mathrm{per}:\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F\rightarrow\mathbb{C}$ is also injective. This suggests that information about the structure of the periods is not also lost by the base extension.
@@
We also extend $\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;w;q)$ $F$-linearity to whole $w\in F\langle F\rangle$.

### v3

@@ 3.2. Explicit formula
$\displaystyle\mathrm{I}_{h}(\omega_{4}^{2m}\omega_{0})$ $\displaystyle=\frac{1}{3^{m}}\frac{1}{(2m)!}\left(\frac{\pi}{3}\right)^{2m}\frac{1}{2}\log 3+(-1)^{-m}\frac{1}{12^{m}}\zeta(2m+1)$ (96)
@@
$\displaystyle=\frac{1}{\sqrt{3}^{k-1}}\frac{1}{(k-1)!}\left(\frac{\pi}{3}\right)^{k-1}\log\frac{\sqrt{3}}{2}$ (107)
$\displaystyle+\frac{1}{\sqrt{-3}^{k-1}}\frac{1}{(k-1)!}\bigg{\{}(-1)^{k}(k-2)!(\mathrm{Li}_{k}(1)+\mathrm{Li}_{k}(-1))$ (108)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\sum_{j=2}^{k}(-1)^{j-1}\frac{(k-1)!}{(k-j)!}\left(\frac{\pi\sqrt{-1}}{3}\right)^{k-j}\bigg{(}\mathrm{Li}_{j}(\xi_{6})+\mathrm{Li}_{j}(-\xi_{6})$ (109)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left.\left.+\frac{1}{(j-1)!}\log 2\left(\frac{\pi\sqrt{-1}}{3}\right)^{j-1}+\frac{2-3j}{2j!}\left(\frac{\pi\sqrt{-1}}{3}\right)^{j}\right)\right\}.$ (110)
@@
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1}{\sqrt{3}^{k-2}}\sum_{j=2\text{: even}}^{k}\frac{(-1)^{\frac{2-j}{2}}}{(k-j)!}\left(\frac{\pi}{3}\right)^{k-j}2^{-j}L(j,\chi_{-3})$ (115)
$\displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\frac{1}{\sqrt{3}^{k-1}}\sum_{j=3\text{: odd}}^{k}\frac{(-1)^{\frac{1-j}{2}}}{(k-j)!}\left(\frac{\pi}{3}\right)^{k-j}2^{-j}(3^{1-j}-1)\zeta(j)$ (116)
and the equations 95 and 98 are shown by setting $k=2m$ and $k=2m+1$. The latter claims are also shown in the same way as the proof of Proposition 3.5. ∎
@@ 4.1. Motivic iterated integrals
Let $\overline{\mathbb{Q}}$ be an algebraic closure of $\mathbb{Q}$ and $F(\subseteq\overline{\mathbb{Q}})$ be a number field. Then, there is a graded commutative Hopf algebra $\mathcal{A}_{F}$ and a graded commutative $\mathcal{A}_{F}$-comodule $\mathcal{H}_{F}$. We investigate the motivic structure of periods in these algebras. Here, we review some properties of them and define motivic iterated integrals.
@@
Let $\mathcal{A}_{F}$ be the fundamental Hopf algebra of the abelian category $\mathrm{MT}(F)$ of mixed Tate motives over $F$. Put $\mathcal{H}_{F}:=\mathcal{A}_{F}\otimes_{\mathbb{Q}}\mathbb{Q}[\tau]$ where $\tau$ is a formal element. Then, $\mathcal{A}_{F}$ and $\mathcal{H}_{F}$ satisfy following conditions.
@@
The period map $\mathrm{per}:\mathcal{H}_{F}\rightarrow\mathbb{C}$ is conjecturally injective (called period conjecture, see [5, p.342, Conjecture 4.118]). This conjecture suggests that information about the structure of the periods is not lost by the motivic lifting.
@@
    $\mathrm{I}^{\mathfrak{a}}(p;a_{1},\dots,a_{k};q):=\rho(\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;a_{1},\dots,a_{k};q))\in\mathcal{A}_{F}^{(k)}.$
@@
    (path connection): For smooth paths $\gamma_{1},\gamma_{2}$ on $\mathbb{C}$ with $\gamma_{1}(1)=\gamma_{2}(0)$,
@@
From here, we investigate the structures of $\mathcal{A}_{F}$ and $\mathcal{H}_{F}$ more precisely. For the $\mathbb{Q}$-subspace $R=\mathcal{O}^{\times}_{F,S}\otimes_{\mathbb{Z}}\mathbb{Q}\subseteq F^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ for some finite set $S$ of finite places of $F$, let $\mathcal{A}_{F,R}$ be a fundamental Hopf algebra of $R$. Then, it can be embedded into $\mathcal{A}_{F}$ (see [8, p.279, Appendix A.4]). We put $\mathcal{H}_{F,R}:=\mathcal{A}_{F,R}\otimes_{\mathbb{Q}}\mathbb{Q}[\tau]$.
@@
Here, $R_{N}:=\mathcal{O}^{\times}_{\mathbb{Q}(\xi_{N}),S_{N}}\otimes_{\mathbb{Z}}\mathbb{Q}\subseteq\mathbb{Q}(\xi_{N})^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ for $S_{N}=\{(1-\xi_{N}^{i})\leavevmode\nobreak\ |\leavevmode\nobreak\ i=1,\dots,N,(i,N)=1\}$.
@@ 4.2. Specific cases
In this subsection, we review some properties of the structure of $\mathcal{H}_{F,R}$ for some specific cases used in Section 5. We consider the two cases: (I) $(F,R)=(\mathbb{Q}(\xi_{4}),\langle 2\rangle_{F_{g}})$, which relates to multiple $L$-values of level 4; (II) $(F,R)=(\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{F_{h}})$, which relates to multiple $L$-values of level 6.
@@
    The $\mathbb{Q}$-linear space $\mathcal{H}_{\mathbb{Q}(\xi_{4}),\langle 2\rangle_{F_{g}}}^{(k)}$ is spanned by motivic iterated integrals formed
@@
    The $\mathbb{Q}$-linear space $\mathcal{H}_{\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{F_{h}}}^{(k)}$ is spanned by motivic iterated integrals formed
@@
    For $k\in\mathbb{Z}_{\geq{1}}$, the elements $\mathrm{I}^{\mathfrak{a}}(0;\delta,\{0\}^{k-1};1)$ $(\delta\in\mu_{4})$ are subject only to the following relations in $\mathcal{A}_{\mathbb{Q}(\xi_{4}),\langle 2\rangle_{F_{g}}}$:
@@
    For $k\in\mathbb{Z}_{\geq{1}}$, the elements $\mathrm{I}^{\mathfrak{a}}(0;\delta,\{0\}^{k-1};1)$ $(\delta\in\mu_{6})$ are subject only to the following relations in $\mathcal{A}_{\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{F_{h}}}$:
@@
$\displaystyle\mathrm{Ker}\Delta^{\prime}\cap\mathcal{H}_{\mathbb{Q}(\xi_{4}),\langle 2\rangle_{F_{g}}}^{(k)}=\mathbb{Q}\tau^{k}\oplus\mathbb{Q}\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{4},\{0\}^{k-1};1);$ (131)
$\displaystyle\mathrm{Ker}\Delta^{\prime}\cap\mathcal{H}_{\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{F_{h}}}^{(k)}=\mathbb{Q}\tau^{k}\oplus\mathbb{Q}\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\xi_{6},\{0\}^{k-1};1).$ (132)
@@ 5.1. Base extension
    The integral path is $\alpha:[0,1]\rightarrow\mathbb{P}^{1}(\mathbb{C})$, $\alpha(t)=1-\sqrt{\frac{1-t}{1+t}}$ and it satisfies $\alpha([0,1])=[0,1]$. Then, each $\mathrm{I}_{g}(\eta_{1}\cdots\eta_{k})$ can be written by a $\mathbb{Q}(\xi_{4})$-linear combination of iterated integrals on $\mathbb{P}^{1}(\mathbb{C})\setminus\{0,1,\pm\sqrt{-1},\infty\}$ having the form
@@
    with $a_{j}\in F_{g}:=\mathbb{Q}(\xi_{4})$ for $j=1,\dots,k$. Also, for each of the terms given by the equation 133, it holds $\widetilde{\mathrm{I}}(a_{i_{1}},a_{i_{2}},a_{i_{3}})\in\langle 2\rangle_{F_{g}}=:R_{g}\subseteq F_{g}^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ if $1\displaystyle 0\leq i_{1}<i_{2}<i_{3}\leq k+$ 1 since $a_{i_{1}},a_{i_{2}},a_{i_{3}}\in\{0,1,\pm\sqrt{-1}\}$.
@@
    $h(X,Y,Z)=X^{2}+XY+Y^{2}-Z^{2}$: for each element of $B_{h}=\left\{\omega_{0}:=\frac{\leavevmode\nobreak\ dx}{x},\omega_{1}:=\frac{\leavevmode\nobreak\ dx}{1-x},\omega_{4}:=\frac{\leavevmode\nobreak\ dx}{x+2y},\omega_{5}:=\frac{\leavevmode\nobreak\ dx}{x(x+2y)},\omega_{6}:=\frac{\leavevmode\nobreak\ dx}{(1-x)(x+2y)}\right\}$, the pullback along $\phi_{h}$ is
@@
    The integral path is $\alpha:[0,1]\rightarrow\mathbb{P}^{1}(\mathbb{C})$, $\alpha(t)=\frac{t-2+\sqrt{4-3t^{2}}}{2(1-t)}$ and it satisfies $[0,1]$. Then, each $\mathrm{I}_{h}(\eta_{1}\cdots\eta_{k})$ can be written by a $\mathbb{Q}(\xi_{6})$-linear combination of iterated integrals on $\mathbb{P}^{1}(\mathbb{C})\setminus\{0,1,-2,\xi_{3}^{\pm 1},\infty\}$ having the form
@@
    with $a_{j}\in F_{h}:=\mathbb{Q}(\xi_{6})$ for $j=1,\dots,k$. Also, for each of the terms given by the equation 139, it holds $\widetilde{\mathrm{I}}(a_{i_{1}},a_{i_{2}},a_{i_{3}})\in\langle 2,3\rangle_{F_{h}}=:R_{h}\subseteq F_{h}^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ if $1\displaystyle 0\leq i_{1}<i_{2}<i_{3}\leq k+$ 1 since $a_{i_{1}},a_{i_{2}},a_{i_{3}}\in\{0,1,-2,\xi_{3}^{\pm 1}\}$.
By the above discussion, for each $f\in\{g,h\}$ and $(\eta_{1},\dots,\eta_{k})\in B_{f}^{(k)}$, $\mathrm{I}_{f}(\eta_{1}\cdots\eta_{k})$ can be written as an $F_{f}$-linear combination of iterated integrals having the form
@@
with $a_{j}\in F_{f}$ for $j=1,\dots,k$. We recall that the $\mathbb{Q}$-linear combination of the equation 140 has a motivic interpretation given by motivic iterated integrals in $\mathcal{H}_{F_{f},R_{f}}$. Then, we consider the base extension of $\mathcal{H}_{F_{f},R_{f}}$ to give motivic interpretations of iterated integrals on $X_{f}$.
@@
For a number field $F\supseteq\mathbb{Q}$ and $R=\mathcal{O}^{\times}_{F,S}\otimes_{\mathbb{Z}}\mathbb{Q}\subseteq F^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$, we extend the following notations to $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$. For $u\otimes c,u_{1}\otimes c_{1},u_{2}\otimes c_{2}\in\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$,
@@
As same as [14, p.19, Proposition 4.4.2], if the period conjecture is true, (i.e. the period map $\mathrm{per}:\mathcal{H}_{F,R}\rightarrow\mathbb{C}$ is injective), then its extension $\mathrm{per}:\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F\rightarrow\mathbb{C}$ is also injective. This suggests that information about the structure of the periods is not preserved by the base extension.
@@
We also extend $\mathrm{I}^{\mathfrak{m}}_{\gamma}(p;w;q)$ $F$-linearity to all $w\in F\langle F\rangle$.